# Improved High-probability Convergence Guarantees of Decentralized SGD

### v2

@@ 2 Problem Setup and Proposed Method
In this section we introduce the problem of decentralized stochastic optimization and the DSGD algorithm. Consider a network of $n\geq 2$ users which can communicate with each other and want to jointly train a model. Formally, the problem can be cast as
$\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}\bigg\{f(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x)\bigg\},$ (2)
@@
where $x\in\mathbb{R}^{d}$ represents model parameters, $f_{i}:\mathbb{R}^{d}\mapsto\mathbb{R}$ is the cost function associated to user $i\in[n]$, given by $f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x;\xi_{i})]$, with $\xi_{i}\in\Xi$ being a random variable governed by some unknown distribution $\mathcal{D}_{i}$, while $\ell:\mathbb{R}^{d}\times\Xi\mapsto\mathbb{R}$ is a loss function. Each user has access to a Stochastic First-order Oracle ($\mathcal{SFO}$), which, when queried by user $i\in[n]$ and input $x\in\mathbb{R}^{d}$, returns the gradient of $\ell$ evaluated at a random sample $\xi_{i}\sim\mathcal{D}_{i}$, i.e., $\nabla\ell(x;\xi_{i})$. The $\mathcal{SFO}$ model subsumes:
  1. 1.
@@
    Batch (offline) learning, where users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a sample $\xi_{i,l}$ uniformly at random, [^4] Note that it is possible to have non-uniform weights, i.e., $f_{i}(x)=\sum_{l\in[m_{i}]}p_{i,l}\ell(x;\xi_{i,l})$, where $p_{i,l}>0$ for all $l\in[m_{i}]$ and $\sum_{l\in[m_{i}]}p_{i,l}=1$, resulting in non-uniform sampling, with sample $\xi_{i,l}$ selected with probability $p_{i,l}$. which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update model parameters;
  2. 2.
@@
    Streaming (i.e., online) learning, where users do not store a local dataset, but in each round observe a random sample $\xi_{i}\sim\mathcal{D}_{i}$ from a data stream, which is used to compute $\nabla\ell(x;\xi_{i})$ and update model parameters.
The communication pattern between users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices representing users, while $E\subset V\times V$ is the set of edges representing communication links between users. [^5] Our work can be readily extended to non-static, time-varying graphs, see the Appendix for a further discussion. To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [diffusion1, diffusion2, diffusion3]. The method consists of the following steps. At the start, users choose a shared step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$ and each user $i\in[n]$ chooses an arbitrary, but deterministically selected initial model $x^{1}_{i}\in\mathbb{R}^{d}$. [^6] While the initial models can be any real vectors, possibly different across users, they needs to be deterministic quantities, for the sake of theoretical analysis. In iteration $t\geq 1$, users query the $\mathcal{SFO}$ with their current model $x^{t}_{i}$ and receive $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$. Users first update their local model via the rule $x_{i}^{t+\nicefrac{{1}}{{2}}}=x^{t}_{i}-\alpha_{t}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$, after which the new model is produced by performing a consensus step, i.e., $x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}x^{t+\nicefrac{{1}}{{2}}}_{j}$, where $\mathcal{N}_{i}\coloneqq\{j\in V:\{i,j\}\in E\}\cup\{i\}$ is the set of users (i.e., neighbours) with whom user $i$ can communicate (including user $i$ itself), while $w_{ij}>0$ is the weight user $i$ assigns to user $j$’s model. The update rule at user $i\in[n]$ can be compactly represented as
$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big).$ (3)
The method is summarized in Algorithm 1. The version of DSGD considered in our work is related to the version based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [nedic-subgrad, consensus+innovation1, consensus+innovation2], whose HP convergence is studied in [dsgd-high-prob].
@@ 3 Main Results
In this section we present the main results. Subsection 3.1 states the preliminaries, Subsection 3.2 provides results for non-convex costs, while Subsection 3.3 presents results for strongly convex costs.
@@ 3.1 Preliminaries
(A1) The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [ran-vr]. Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix. It can be shown that (A1) implies $\lambda\in[0,1)$, see [Horn_Johnson_2012].
(A2) The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
(A3) Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [ghadimi2013stochastic, pmlr-v119-koloskova20a, ran-improved].
(A4) The stochastic quantities satisfy the following:
@@
    The stochastic gradients are unbiased, i.e., for any $i\in[n]$, $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
    $\mathbb{E}[\nabla\ell(x;\xi_{i}^{t})\>|\>\mathcal{F}_{t}]=\nabla f_{i}(x).$
@@
    The noise at each user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian, i.e., for all $t\geq 1$ and any $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
@@
The first condition in (A4) is standard in decentralized stochastic optimization [ran-improved, yu2025decentralized], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [nemirovski2009robust, ghadimi2013stochastic, li-orabona, li2020high, pmlr-v99-harvey19a, liu2023high, pmlr-v206-bajovic23a]. [^7] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [madden2020high] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
(A5) The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^8] With at least one of $A$ or $B$ being strictly positive. we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [NIPS2017_f7552665, dist-learn-noncvx, pmlr-v119-koloskova20a, wang-cooperative]. Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [dsgd-high-prob], as it allows users’ gradients to grow with the global gradient. Compared to [pmlr-v119-koloskova20a], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition. While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [pmlr-v119-koloskova20a], to keep the proofs simple and instructive, we will use assumption (A5).
(A6) Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
Assumption (A6) is used in the analysis of strongly convex costs. It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a, wang-cooperative] which is in general not required for centralized SGD. The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [ran-improved], however, this is beyond the scope of the current work.
Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$ denote the network-wise average model. Note that $\overline{x}^{t}$ is an ideal model, used for analysis purposes. Using (3) and the fact that $W$ is doubly stochastic, it can be readily seen that
$\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t},$
where $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ is the average stochastic gradient. Let $\overline{z}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}z_{i}^{t}$ denote the average noise. We then have the following result.
@@
Lemma 1 provides an important deterministic descent-type inequality for DSGD and is the starting point for our analysis. The right-hand side of the above inequality consists of terms that arise in centralized SGD plus a consensus gap term $\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$, which stems from the decentralized nature of the algorithm and bounding this term is crucial to ensure convergence. To finish this subsection, we provide a technical result on the behaviour of the stochastic noise in the HP sense. Let $\sigma^{2}\coloneqq\frac{1}{n}\sum_{i\in[n]}\sigma^{2}_{i}$ be the average noise parameter. We then have the following result.
@@
Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections. We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
@@ 3.2 Non-convex Costs
In this subsection we present results for non-convex costs. Let $\Delta_{x}\coloneqq\frac{1}{n}\sum_{i\in[n]}\|x^{1}_{i}-\overline{x}^{1}\|^{2}$ denote the initial consensus gap. We then have the following result.
@@
Let (A1) and (A5) hold. We then have, for any $t\geq 1$
Lemma 3 provides a deterministic bound on the consensus gap and is used to bound the resulting MGF. Define $\Delta_{f}\coloneqq f(\overline{x}^{1})-f^{\star}$ and $\overline{\sigma}\coloneqq\max_{i\in[n]}\sigma_{i}$. Building on Lemmas 1 and 3, we get the following result.
@@
Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few perspectives.
@@
Centralized SGD and linear speed-up. Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [liu2023high], we can see that the rate from Theorem 1 similarly consists of two terms of different orders, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$. Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [pmlr-v119-koloskova20a], and shows that DSGD maintains the benefits of decentralized learning, in the HP sense.
@@
Network effect and transient time. We can see from the bound in Theorem 1 that network connectivity only affects the higher-order terms decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is again consistent with MSE results, e.g., [pmlr-v119-koloskova20a]. Moreover, we can use the rate in Theorem 1 to establish the transient time of DSGD in the HP sense, which can be shown to be of the order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD stemming from MSE results, see, e.g., [unified-refined, Table I]. For details on the derivation of the transient time, see the Appendix.
@@
Effects of noise and heterogeneity. The noise (through $\sigma,\overline{\sigma}$) and heterogeneity (through $A,B$) mainly affect the higher-order terms, [^9] The effect of noise on the leading term can be eliminated by choosing the step-size $\alpha=\min\Big\{C,\frac{\sqrt{n}}{\sqrt{dLT}}\Big\}$. which is again consistent with MSE results. Note that, if $A=0$, the bound in Theorem 1 recovers the convergence rate of noiseless gradient descent, in the sense that, when $\sigma=A=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
@@
Comparison with [dsgd-high-prob]. The authors in [dsgd-high-prob] study HP guarantees of DSGD over undirected time-varying networks, establishing a rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$, using a time-varying step-size schedule $\alpha_{t}=\frac{C}{\sqrt{t+1}}$, for some $C>0$ and assuming uniformly bounded gradients. As we discuss next, our results provide several important improvements. First, we remove the uniformly bounded gradient assumption and replace it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details. Next, the authors in [dsgd-high-prob] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise. In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, as all terms converge at the same $\mathcal{O}\Big(\frac{1}{\sqrt{T}}\Big)$ rate. However, it would be impossible to obtain a linear speed-up in [dsgd-high-prob] even under a fixed step-size, as the authors show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity. On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used (see the Appendix for a detailed discussion on the results needed to achieve linear speed-up). [^10] This is also consistent with existing MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where linear speed-up for non-convex costs is achieved using a fixed step-size. Finally, while we consider static networks, our results can be readily extended to the time-varying setting considered in [dsgd-high-prob], see the Appendix for details.

### v3

@@ 2 Problem Setup and DSGD
In this section we introduce the problem of interest and the DSGD algorithm. Consider a network of $n\geq 2$ users who can communicate with each other and want to jointly train a model. Formally, the problem can be cast as
$\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}\Big\{f(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x)\Big\},$ (2)
@@
where $x\in\mathbb{R}^{d}$ represents model parameters, $f_{i}:\mathbb{R}^{d}\mapsto\mathbb{R}$ is the cost function of user $i\in[n]$, given by $f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x;\xi_{i})]$, with $\xi_{i}\in\Xi$ being a random variable governed by an unknown distribution $\mathcal{D}_{i}$, while $\ell:\mathbb{R}^{d}\times\Xi\mapsto\mathbb{R}$ is a loss function. Each user has access to a Stochastic First-order Oracle ($\mathcal{SFO}$), which, when queried by user $i\in[n]$ with input $x\in\mathbb{R}^{d}$, returns the gradient of $\ell$ evaluated at a random sample, i.e., $\nabla\ell(x;\xi_{i})$. The $\mathcal{SFO}$ model subsumes:
1. Batch (offline) learning
@@
- users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a random sample $\xi_{i,l}$, which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update the model;
2. Streaming (online) learning
@@
- users do not store a local dataset, but in each round observe a random sample $\xi_{i}$, which is used to compute $\nabla\ell(x;\xi_{i})$ and update the model.
@@
The communication pattern between users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices representing users, while $E\subset V\times V$ is the set of edges representing communication links between users. To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [42, 13, 15]. The method consists of the following steps. First, users choose a shared step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$ and each user $i\in[n]$ chooses an arbitrary, but deterministically selected initial model $x^{1}_{i}\in\mathbb{R}^{d}$. [^4] While initial models can be any real vectors, possibly different across users, for analysis purposes they need to be deterministic. In iteration $t\geq 1$, users query the $\mathcal{SFO}$ with their current model $x^{t}_{i}$ and receive $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$. Users then update their models via the rule
$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big),$ (3)
which consists of a local update, followed by a consensus step, where $\mathcal{N}_{i}\coloneqq\{j\in V:\{i,j\}\in E\}\cup\{i\}$ is the set of users (i.e., neighbours) with whom user $i$ can communicate (including $i$ itself), while $w_{ij}>0$ is the weight user $i$ assigns to user $j$’s model. The method is summarized in Algorithm 1. The version of DSGD considered in our work is closely related to the variant based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [53, 28, 29], whose HP convergence is studied by [44].
@@ 3 Main Results
In this section we present the main results. Subsection 3.1 states the preliminaries, while Subsections 3.2 and 3.3 provide results for non-convex and strongly convex costs.
@@ 3.1 Preliminaries
The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [74]. [^5] Our work can be readily extended to the setting of time-varying networks, see Appendix J for a detailed discussion. Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix. It can be shown that (A1) implies $\lambda\in[0,1)$, see, e.g., [22].
The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [20, 75].
The stochastic quantities satisfy the following:
@@
    The stochastic gradients are unbiased, i.e., for any $i\in[n]$, $t\geq 1$ and $\mathcal{F}_{t}$-measurable vector $x\in\mathbb{R}^{d}$, we have $\mathbb{E}[\nabla\ell(x;\xi_{i}^{t})\>|\>\mathcal{F}_{t}]=\nabla f_{i}(x)$.
@@
    The noise at user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian (light-tailed), i.e., for all $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
@@
The first condition in (A4) is standard in decentralized stochastic optimization [75, 79], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form in (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [54, 20, 36, 37, 21, 39, 5]. [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [46] and references therein, the focus of this work is on sub-Gaussian noise.
The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
A heterogeneity bound of the type in (A5) is in fact required to ensure that DSGD converges for non-convex costs [38, 14, 33, 72]. Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [44], as it allows users’ gradients to grow with the global gradient. Compared to [33], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition. While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [33], to keep the proofs simple and instructive, we will use assumption (A5).
Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
Assumption (A6) is used for the strongly convex case. It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [24, 33, 72]. The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [75], however, this is beyond the scope of the current work. Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$, $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ and $\overline{z}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}z_{i}^{t}$ denote the network-average model, stochastic gradient and noise, respectively. Using (3) and the double stochasticity of $W$, it follows that
$\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t}.$
We then have the following result.
@@
Lemma 1 provides an important deterministic descent-type inequality for DSGD and is the starting point for our analysis. The right-hand side of the above inequality consists of terms that arise in centralized SGD plus a consensus gap term $\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$, which stems from the decentralized nature of the algorithm and bounding this term is crucial to ensure convergence. To close the subsection, we provide an important result on the behaviour of the stochastic noise in the HP sense. Let $\sigma^{2}\coloneqq\frac{1}{n}\sum_{i\in[n]}\sigma^{2}_{i}$ be the average noise parameter. We then have the following result.
@@
Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections. Note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and Appendix K.
@@ 3.2 Non-Convex Costs
In this subsection we present results for non-convex costs. Let $\Delta_{x}\coloneqq\frac{1}{n}\sum_{i\in[n]}\|x^{1}_{i}-\overline{x}^{1}\|^{2}$ and $\Delta_{f}\coloneqq f(\overline{x}^{1})-f^{\star}$ denote the initial consensus and optimality gap.
@@
If (A1) and (A5) hold, then for any $t\geq 1$
$\frac{1}{n}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\leq 2\lambda^{2t}\Delta_{x}+\frac{4\lambda^{2}A^{2}}{1-\lambda}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\Big[\|z^{k}_{i}\|^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\Big].$
Lemma 3 provides a useful deterministic bound on the consensus gap. Define $\overline{\sigma}\coloneqq\max_{i\in[n]}\sigma_{i}$. Building on Lemmas 1 and 3, we get the following result.
@@
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
then for any $\delta\in(0,1)$, with probability at least $1-\delta$
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
@@
Theorem 1 establishes HP convergence of DSGD for general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few different perspectives.
Centralized SGD and linear speed-up.
@@
Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [39], we can see that the rate in Theorem 1 also consists of two different order terms, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$. Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [33], showing that DSGD retains the benefits of decentralized learning in the HP sense.
Network effect and transient time.
@@
We can see from the bound in Theorem 1 that network connectivity only affects higher-order terms, decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is consistent with MSE results, e.g., [33]. Moreover, the rate in Theorem 1 implies that the transient time of DSGD in the HP sense is of order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD implied by MSE rates, see, e.g., Table I in [1]. For details on the derivation of the transient time, see Appendix H.
Effects of noise and heterogeneity.
@@
The noise (through $\sigma,\overline{\sigma}$) and heterogeneity (through $A,B$) mainly affect the higher-order terms, [^7] The effect of noise on the leading term can be eliminated by choosing the step-size $\alpha=\min\Big\{C,\frac{\sqrt{n}}{\sqrt{dLT}}\Big\}$. which is again consistent with MSE results. Note that, if $A=0$, the bound in Theorem 1 recovers the convergence rate of noiseless gradient descent, in the sense that, when $\sigma=A=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
Comparison with [44].
@@
The authors in [44] study HP guarantees of DSGD over undirected time-varying networks, under a time-varying step-size schedule $\alpha_{t}=\frac{a}{\sqrt{t+1}}$, assuming uniformly bounded gradients and establish the rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$. As highlighted next, we provide several important improvements. First, we remove the uniformly bounded gradient assumption, replacing it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details. Next, [44] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise. In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix E for a detailed discussion. However, it would be impossible for [44] to obtain a linear speed-up even under a fixed step-size, as they show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity. On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix I for detailed discussion on the results needed to achieve linear speed-up. [^8] This is consistent with MSE results, e.g., [33, 75, 1], where linear speed-up in non-convex case is achieved only via a fixed step-size. Finally, while we consider static networks, our results can be extended to the time-varying setting in [44], see Appendix J for details.